A theater group made appearances in two cities. The hotel charge before tax in the second city was
$500
lower than in the first. The tax in the first city was
4%
, and the tax in the second city was
8%
. The total hotel tax paid for the two cities was
$380
. How much was the hotel charge in each city before tax?

Answer by jorel555(1290)

m+n=380
.08m-.04n=500
2m+n=12500
m=12500-380=$12120 was the charge in the SECOND city; $11740 was the charge in the first!!!!!!!!!!!!